OpenStudy (anonymous):

Help please :/ Find the following infinite sum, if it exists. (LOOK AT RESPONSE)

OpenStudy (anonymous):

\[\sum_{n-1}^{\infty}8\left(\begin{matrix}4 \\ 3\end{matrix}\right)^{n-1}\]

OpenStudy (anonymous):

I can rewrite it as: \[8\sum _{n=0}^{\infty } \left(\frac{4}{3}\right)^n\]

OpenStudy (anonymous):

4/3 is bigger than 1 thus, the sum does not converge